I need help with subtracting radicals PLEASE!!!!!!
The problem is 5sqrt72-3sqrt48-4sqrt128=

Answer by venugopalramana(3286)

5sqrt72-3sqrt48-4sqrt128=?...the procedure is to factorise the radicals and take as much as possible out side the radical and simplify as per the classification
for ex...sqrt72=sqrt2*2*2*3*3..there are three 2's and two 3's .so we can take two each of them out side the radical of square root,leaving only one 2 in the radical...
sqrt72=sqrt2*2*2*3*3=2*3sqrt2=6sqrt2...similarly
sqrt48=sqrt2*2*2*2*3=2*2sqrt3=4sqrt3
sqrt128=sqrt2*2*2*2*2*2*2=2*2*2sqrt2=8sqrt2..
now you should know that in their simplest form like sqrt2,sqrt3,sqrt5...etc can not be simplified any further and can be added or subtracted within their category only..
so we have now 5sqrt72-3sqrt48-4sqrt128=5*6sqrt2-3*4sqrt3-4*8sqrt2=sqrt2(30-32)-12sqrt3=-2sqrt2-12sqrt3...answer